# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

### v2

@@ 4.3 Constrained Objective for Non-Privileged Labels ($j\in\bar{\mathcal{P}}$)
$\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|\exists t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},j)\text{ s.t. }j\in\bar{\mathcal{P}}}\left[\max(0,\ell(\mathbf{w}_{j};x_{i},y_{ij})-\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})-\epsilon)\right].$ (17)
@@ 4.4 Group Robust Optimization Formulation
[…]
Here, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are adaptive weights representing the importance assigned to each group’s loss. The inner maximization finds the worst-case distribution over group losses (by increasing the weight $\alpha$ for the group with higher current loss), while the outer minimization seeks model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}$ that perform well even under this worst-case weighting. This formulation inherently promotes robustness and fairness by preventing the optimization from disproportionately favouring one group at the expense of the other.
@@ 4.5 Optimization Algorithm
The group weights $\alpha^{(s+1)}$ are updated using a mirror ascent step. A crucial aspect for the stability and effectiveness of this step is loss scaling. Instead of directly using the raw current loss values $\mathcal{L}_{g}^{(s)}$ (for $g\in\{\mathcal{P},\bar{\mathcal{P}}\}$) in the exponent, we use a scaled version. Specifically, for each group $g$, we maintain a running average of its loss from previous steps, denoted ¯$\bar{\mathcal{L}}_{g}^{\text{avg}}$. The term used to update the weights is then based on the relative change of the current loss from this average, for example, a scaled difference like $(\mathcal{L}_{g}^{(s)}-\bar{\mathcal{L}}_{g}^{\text{avg}})/(\bar{\mathcal{L}}_{g}^{\text{avg}}+\delta)$, where $\delta$ is a small constant for numerical stability. This normalization prevents instability due to scale disparities in group loss values and makes the optimization more sensitive to relative improvements or regressions, rather than absolute magnitudes.
@@
1:Initialize: Model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}^{(0)}$ (e.g., from supervised fine-tuning), set group weights $\alpha_{\mathcal{P}}^{(0)},\alpha_{\bar{\mathcal{P}}}^{(0)}$
2:Input: Dataset $\mathcal{D}$, reference parameters $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$, hyperparameters $\beta,\epsilon,\eta_{\mathbf{w}},\eta_{\alpha}$.
@@
10:      Update model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}^{(s+1)}$ via mirror descent using weighted gradients.
@@
12:Return Optimized parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}^{(S)}$.
@@ 5 Experimental Setup
Datasets and Preprocessing: We evaluate FairPO on two standard multi-label image classification benchmarks: MS-COCO 2014222MS-COCO: https://cocodataset.org/ (80 object categories) (Lin et al., 2015) and NUS-WIDE [^3] NUS-WIDE: https://github.com/NExTplusplus/NUS-WIDE/ (81 concept labels) (Chua et al., 2009). For partitioning, we designate the 20% least frequent labels in the training set of each dataset as the privileged group ($\mathcal{P}$), with the remaining 80% forming the non-privileged group ($\bar{\mathcal{P}}$). We use standard train/validation/test splits for both datasets (see Appendix C for more details).
Model and Implementation Details: Our base model is a Vision Transformer (ViT) (Dosovitskiy et al., 2021) pretrained on ImageNet [^4] ViT Model: https://huggingface.co/docs/transformers/en/model_doc/vit. The FairPO framework, including the classifier head and label-specific parameters $\mathbf{w}_{t}$, is fine-tuned on top of the frozen ViT features. The reference parameters $\hat{\mathbf{w}}$ for FairPO are obtained by standard supervised fine-tuning (SFT) of the same ViT architecture with a BCE loss on the target MLC dataset. Hyperparameters like learning rates ($\eta_{\mathbf{w}},\eta_{\alpha}$), DPO’s $\beta$, SimPO’s $\gamma$, CPO’s $\lambda_{\text{CPO}}$, and the non-privileged slack $\epsilon$ are tuned via a grid search on the validation set (see Appendix F for more details).
Evaluation Metrics: Model performance is assessed using established multi-label metrics, calculated separately for $\mathcal{P}$ and $\bar{\mathcal{P}}$ label sets. These include Mean Average Precision (mAP), which measures the average precision across all labels; Sample F1 Score (F1-Sample), the harmonic mean of precision and recall averaged per sample; Exact Match Ratio (EMR), the proportion of samples with perfectly predicted label sets; and Multi-label Accuracy (Accuracy), representing element-wise prediction correctness averaged over all labels and instances.
Baselines: FairPO is compared against: (1) BCE-SFT, standard per-label Binary Cross-Entropy (BCE) fine-tuning, which also provides reference parameters $\hat{\mathbf{w}}$. (2) BCE-SFT + Privileged Re-Weighting (RW), which statically increases loss weights for the privileged group $\mathcal{P}$. (3) Group DRO + BCE (Sagawa et al., 2019), applying Group DRO to minimize the worst-group BCE loss between $\mathcal{P}$ and $\bar{\mathcal{P}}$ (see Appendix D for more details).
FairPO Variants: We evaluate three main variants of our FairPO framework, each differing in the specific preference loss formulation used for the privileged group $\mathcal{P}$ (as detailed in Section 4.2): FairPO-DPO utilizes the DPO-inspired conditional preference loss (Eq. 11). FairPO-SimPO employs a SimPO-inspired (Meng et al., 2024) reference-free loss, adapting the preference term from Eq. 12 within the conditional structure of Eq. 11. Lastly, FairPO-CPO uses a CPO-inspired (Xu et al., 2024) reference-free loss combined with an NLL-like regularization component, as formulated in Eq. 14 (see Appendix E for more details).
@@ 6 Results and Analysis
Overview and Performance Summary: We evaluate FairPO against baselines on MS-COCO (Lin et al., 2015) and NUS-WIDE (Chua et al., 2009). Tables 1 and 2 detail performance for privileged ($\mathcal{P}$, 20% least frequent) and non-privileged ($\bar{\mathcal{P}}$, 80% most frequent) label sets. The $\Delta\text{mAP}(\mathcal{P})$ column highlights mAP gains for $\mathcal{P}$ labels over the BCE SFT baseline. A consistent trend emerges: FairPO variants, particularly FairPO-CPO, significantly improve $\mathcal{P}$ set performance while robustly maintaining $\bar{\mathcal{P}}$ set metrics. This addresses the common trade-off where gains in one group detrimentally affect the other, a challenge often faced by simpler baselines.
@@
Comparative Analysis of FairPO Variants: The FairPO variants differ in their preference mechanisms: FairPO-DPO utilizes a reference model $\hat{\mathbf{w}}$ in its log-ratio preference term. This relative formulation can stabilize training by anchoring scores to a pre-trained baseline, making the preference signal robust to the absolute scale of model outputs. This likely contributes to its strong EMR-$\mathcal{P}$ gains, indicating better resolution of complex multi-label confusion. FairPO-CPO, often the top performer, employs a reference-free contrastive term $\log(m(x_{i};\text{preferred})/m(x_{i};\text{dispreferred}))$. Its strength is amplified by an NLL-like regularizer (realized as $\ell_{\text{BCE}}$ for all privileged labels in Eq. 14). This dual focus on relative preference and absolute correctness for $\mathcal{P}$ underpins its superior $\mathcal{P}$ metrics while maintaining robust $\bar{\mathcal{P}}$ performance. FairPO-SimPO, another reference-free method using a fixed margin $\gamma$, generally shows improvements but lags behind DPO and CPO variants. This suggests that a fixed margin without DPO’s reference anchoring or CPO’s explicit NLL-like regularization might be less robust or harder to tune optimally for diverse preference pairs in MLC.
Impact on Exact Match Ratio: The substantial improvements in EMR-$\mathcal{P}$ by FairPO-DPO and FairPO-CPO are particularly significant. EMR, requiring perfect prediction for the $\mathcal{P}$ label subset, is highly sensitive to confusion. FairPO’s preference optimization, by directly addressing ambiguities between true labels and their specific confusing counterparts, enhances discriminative power for these challenging cases. This leads to more accurate complete label set predictions for the privileged group. The GRPO mechanism ensures these targeted gains do not disproportionately penalize the $\bar{\mathcal{P}}$ group, demonstrating FairPO’s capacity to learn a fairer and more robust multi-label classifier.
@@ 7 Ablation Studies
We conduct ablation studies on MS-COCO using FairPO-CPO to assess component contributions. Results are in Table 3 and Table 4.
@@
Method Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 48.51 35.30 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) (-16.92) (+0.07) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 63.12 31.95 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) (-2.31) (-3.28) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 56.70 34.15 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) (-8.73) (-1.08)
Table 3 shows that removing the preference loss for $\mathcal{P}$ (using BCE instead) significantly reduces $\mathcal{P}$ gains, underscoring its criticality. Removing the $\bar{\mathcal{P}}$ constraint causes a notable drop in $\bar{\mathcal{P}}$ performance, affirming its protective role. Using fixed weights instead of GRPO degrades performance for both groups, highlighting GRPO’s superior adaptive balancing.
@@
Table 4 details the impact of specific choices within the $\mathcal{P}$ loss formulation. Restricting preference learning to only Confusing Negatives (when $y_{il}=+1$) severely curtails its effectiveness. Given that positive instances of privileged labels can be extremely rare, this makes the preference signal exceptionally sparse. This approach neglects to train the model on ranking true negatives ($y_{il}=0$) against confusing positives—a scenario that forms the vast majority of instances for these labels. This omission of crucial learning signals for most situations involving $\mathcal{P}$ labels results in a profound degradation in performance for this set. Omitting the BCE Fallback for non-confusing $\mathcal{P}$ instances (i.e., no loss if $S_{il}=\emptyset$) also clearly degrades performance. This standard classification signal for easier cases is vital for model stability and reinforcing correct predictions within $\mathcal{P}$ when direct preference learning is not active. Both aspects of our full preference formulation are thus shown to be essential. This demonstrates that preference loss alone is insufficient for stable training, and standard losses play a critical complementary role even under targeted preference optimization.
@@ 8 Discussion
Conclusion: We introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in multi-label classification. Our experiments, particularly with the FairPO-CPO variant, underscore the efficacy of this approach in navigating complex fairness-performance trade-offs, highlighting the value of nuanced preference signals for challenging discrimination tasks.
Limitations: Despite its promising results, FairPO presents technical limitations. The dynamic construction of the ‘confusing set’ $S_{il}$ based on evolving model scores can lead to optimization instability or sparse preference signals, especially for rare/easily-separated privileged labels. Furthermore, the efficacy of FairPO-DPO and the non-privileged constraint hinges on a well-calibrated reference model $\hat{\mathbf{w}}$; its deficiencies can distort DPO log-ratios or misguide the constraint. Finally, GRPO’s adaptive balancing of heterogeneous loss terms ($\mathcal{L}_{\mathcal{P}}$ combining preference and BCE, $\mathcal{L}_{\bar{\mathcal{P}}}$ with a hinge loss) may require careful tuning of $\eta_{\alpha}$ and loss scaling for stable convergence, the theoretical properties of which remain an open area. While these limitations highlight the complexity of fair MLC optimization, they also suggest promising directions for refinement and extension, which we talk about next.
Future Work: Future work will focus on several key directions. A primary goal is to extend FairPO’s principles to multi-label attribute generation (see Appendix A for a conceptual outline), aiming to foster fair and diverse outputs. We also plan comprehensive empirical validation across a wider array of datasets and modalities, including text, to assess generalizability. Further research will involve in-depth analysis of different label partitioning strategies beyond frequency, and pursuing theoretical insights into convergence properties for FairPO.
Acknowledgments and Disclosure of Funding
The authors would like to express their gratitude to the members of the CSE Department and CMInDS at IIT Bombay for their support and for providing a conducive research environment. We also thank the anonymous reviewers for their valuable feedback which helped improve the quality of this manuscript. We acknowledge the use of high-performance computing resources that were instrumental in conducting the experiments presented in this work.
Funding Disclosure This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors for the submitted work, beyond the general support and infrastructure provided by IIT Bombay to its faculty and students.
Competing Interests The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v3

@@ 3.3 Constrained Objective for Non-Privileged Labels ($j\in\bar{\mathcal{P}}$)
$\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},j)\text{ s.t. }j\in\bar{\mathcal{P}}}\left[\max(0,\ell(\mathbf{w}_{j};x_{i},y_{ij})-\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})-\epsilon)\right].$ (17)
@@ 3.4 Group Robust Optimization Formulation
$\min_{\{\mathbf{w}_{t}|t\in\mathcal{T}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\,\geq\,0,\,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\,=\,1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})\right].$ (18)
Here, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are adaptive weights representing the importance assigned to each group’s loss. The inner maximization finds the worst-case distribution over group losses (by increasing the weight $\alpha$ for the group with higher current loss), while the outer minimization seeks model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}$ that perform well even under this worst-case weighting. This formulation inherently promotes robustness and fairness by preventing the optimization from disproportionately favouring one group at the expense of the other.
@@ 3.5 Optimization Algorithm
The group weights $\alpha^{(s+1)}$ are updated using a mirror ascent step. A crucial aspect for the stability and effectiveness of this step is loss scaling. Instead of directly using the raw current loss values $\mathcal{L}_{g}^{(s)}$ (for $g\in\{\mathcal{P},\bar{\mathcal{P}}\}$) in the exponent, we use a scaled version. Specifically, for each group $g$, we maintain a running average of its loss from previous steps, denoted $\bar{\mathcal{L}}_{g}^{\text{avg}}$. The term used to update the weights is then based on the relative change of the current loss from this average, for example, a scaled difference like $(\mathcal{L}_{g}^{(s)}-\bar{\mathcal{L}}_{g}^{\text{avg}})/(\bar{\mathcal{L}}_{g}^{\text{avg}}+\delta)$, where $\delta$ is a small constant for numerical stability. This normalization prevents instability due to scale disparities in group loss values and makes the optimization more sensitive to relative improvements or regressions, rather than absolute magnitudes.
@@
1:Initialize: Model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(0)}$ (e.g., from supervised fine-tuning), set group weights $\alpha_{\mathcal{P}}^{(0)},\alpha_{\bar{\mathcal{P}}}^{(0)}$
2:Input: Dataset $\mathcal{D}$, reference parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$, hyperparameters $\beta,\epsilon,\eta_{\mathbf{w}},\eta_{\alpha}$.
@@
10:    Update model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(s+1)}$ via mirror descent using weighted gradients.
@@
12:Return Optimized parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(S)}$.
@@ 4 Experimental Setup
We evaluate FairPO on two standard multi-label image classification benchmarks: MS-COCO 2014 (Lin et al., 2015) and NUS-WIDE (Chua et al., 2009). For our fairness setup, we define the privileged group ($\mathcal{P}$) as the 20% least frequent labels in each training set, with the remaining 80% forming the non-privileged group ($\bar{\mathcal{P}}$). Our base model is a Vision Transformer (ViT) (Dosovitskiy et al., 2021), where we fine-tune label-specific classifier heads on top of its features. The reference model parameters $\hat{\mathbf{w}}$ are obtained from a standard supervised fine-tuning (SFT) of this architecture with a Binary Cross-Entropy (BCE) loss.
Performance is assessed separately for $\mathcal{P}$ and $\bar{\mathcal{P}}$ sets using standard metrics: Mean Average Precision (mAP), Sample F1, Exact Match Ratio (EMR), and Accuracy. We compare FairPO against three baselines: (1) BCE-SFT, which also serves as our reference model; (2) BCE-SFT + Re-Weighting (RW), a simple loss up-weighting for $\mathcal{P}$ labels; and (3) Group DRO + BCE (Sagawa et al., 2020), which minimizes the worst-group loss. We test three variants of our framework, differing in their preference loss for privileged labels: FairPO-DPO (Eq. 11), FairPO-SimPO (Eq. 12), and FairPO-CPO (Eq. 14). Comprehensive details on datasets, preprocessing, model architecture, baseline implementations, and hyperparameter tuning are deferred to Appendices C, D, and E.
@@ 5 Results and Analysis
Tables 1 and 2 detail our findings. The standard BCE-SFT baseline confirms the fairness problem central to our motivation: on MS-COCO, a significant performance gap exists between the privileged ($\mathcal{P}$) labels (86.32% mAP) and the non-privileged ($\bar{\mathcal{P}}$) labels (90.65% mAP). This disparity highlights how standard aggregate losses can fail to ensure equitable performance for rare or difficult categories.
Our results demonstrate that FairPO variants, particularly FairPO-CPO, effectively address this challenge by navigating the fairness-performance trade-off in a controlled manner. On MS-COCO, FairPO-CPO achieves a substantial +3.44% mAP gain for $\mathcal{P}$ labels, while performance on $\bar{\mathcal{P}}$ labels dips by a negligible 0.31%. Similarly, on NUS-WIDE, it yields a +3.59% mAP gain for a minor 0.41% drop. This is not an uncontrolled side effect but a managed outcome, enforced by the GRPO mechanism and the constrained loss (Eq. 17), which together prevent significant harm to the non-privileged group while targeting improvements for the privileged one. By directly optimizing the model to rank true labels higher than their specific, dynamically identified confusing counterparts, FairPO sharpens discriminative power for the most challenging cases.
Among the variants, the reference-free FairPO-CPO proves most effective. This directly addresses a potential limitation of DPO-style methods, which can be sensitive to the quality of the reference model. FairPO-CPO’s strong performance, likely due to its dual objective of optimizing both relative preference and absolute correctness (via its NLL-like regularizer), demonstrates its robustness and versatility, making it a more practical choice than the reference-dependent FairPO-DPO or the fixed-margin FairPO-SimPO.
@@ 6 Ablation Studies
Method Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 48.51 35.30 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) (-16.92) (+0.07) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 63.12 31.95 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) (-2.31) (-3.28) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 56.70 34.15 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) (-8.73) (-1.08) Global CPO (on all labels) 88.55 90.68 62.75 64.85 96.95 98.11 55.20 37.15 +2.23 (No $\mathcal{P}$$\bar{\mathcal{P}}$ split or GRPO) (-1.21) (+0.34) (-1.26) (-0.47) (-1.08) (+0.05) (-10.23) (+1.92)
@@
We conduct ablation studies on MS-COCO using FairPO-CPO to assess component contributions (Tables 3 and 4). Our analysis of core components (Table 3) confirms the criticality of each part. Replacing the preference loss with BCE reduces privileged gains, removing the $\bar{\mathcal{P}}$ constraint harms non-privileged performance, and using fixed weights instead of GRPO degrades both groups. To further isolate the value of our fairness framework, we tested a Global CPO variant that applies the preference objective to all labels without the privileged/non-privileged split or GRPO. While this proves the preference objective is a strong general-purpose loss (achieving a +2.23% mAP gain over BCE), the full FairPO framework is decisively superior for the targeted fairness task. By applying CPO globally, the performance gains are unfocused, resulting in a 1.21 point lower mAP on the privileged set. This demonstrates that the full FairPO architecture is crucial for focusing the model’s capacity to resolve the hardest discrimination challenges within the privileged group.
Furthermore, the design of the preference objective itself is crucial (Table 4). Restricting it to only Confusing Negatives (for $y_{il}=+1$) causes a profound performance drop, as this neglects the far more common scenario of ranking true negatives below confusing positives for rare labels, making the learning signal exceptionally sparse. Similarly, removing the BCE Fallback for non-confusing instances degrades performance, proving that a standard classification signal on easier cases is vital for model stability. Both aspects of our full formulation are thus essential, demonstrating that while targeted preference optimization is powerful, it must be complemented by standard losses and robust balancing for stable, effective training.
7 Related Work
Recent efforts in fair MLC address complex challenges like label imbalance impacting tail labels (Guo and Wang, 2021), subjective fairness (Liu et al., 2023), and class-incremental learning (Dong et al., 2025). FairPO contributes a novel approach by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets. It applies distinct, fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and manages the trade-off with a robustness framework, differentiating our targeted approach from prior work.
We adapt recent advances in preference optimization, originally developed for aligning LLMs (Ouyang and Others, 2022; Christiano et al., 2023). Techniques like Direct Preference Optimization (DPO) (Rafailov et al., 2024) and its reference-free variants CPO (Xu et al., 2024) and SimPO (Meng et al., 2024) optimize policies from preference pairs. Rather than ranking entire outputs, we repurpose these methods to specifically differentiate true label scores from their dynamically identified confusing counterparts within the privileged set, thereby sharpening critical decision boundaries. To balance our objectives, we employ a Group Robust Optimization strategy inspired by Group DRO (Sagawa et al., 2020; Rice et al., 2021) and GRPO (Ramesh et al., 2024). While these methods typically balance performance across data or preference groups, FairPO uniquely defines its groups by our label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$). It then uses GRPO’s adaptive weighting to balance their distinct, custom-formulated loss objectives, providing a principled mechanism for managing the specific fairness-performance trade-offs in our MLC context.
@@ 8 Discussion
In conclusion, we introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in multi-label classification. Our experiments, particularly with the FairPO-CPO variant, highlight the value of nuanced preference signals for navigating complex fairness-performance trade-offs in challenging discrimination tasks. While promising, FairPO has technical limitations: its dynamic confusing set can lead to instability or sparse signals for rare labels; the DPO-based variant and non-privileged constraint rely on a well-calibrated reference model; and GRPO’s balancing of heterogeneous losses requires careful tuning and lacks theoretical convergence guarantees. These challenges also define our future work, which includes extending FairPO’s principles to multi-label attribute generation (Appendix A), conducting comprehensive empirical validation on more datasets and modalities, analyzing alternative label partitioning strategies, and pursuing theoretical insights into the framework’s convergence properties.